OpenStudy (rushda99):

Please help... For complex numbers z satisfying arg( z − u) = π/4, find the least possible value of |z|. u = -3i Please show me how to solve this GEOMETRICALLY (with a half-line).

OpenStudy (phi):

z- u= a+bi - (-3i)= a+(b+3) i and tan (b+3)/a = 1 I would start with that, and start drawing figures

OpenStudy (rushda99):

I marked the point u at (0,-3) then drew a line from that point at a 45 deg angle to the horizontal axis. Now what do I do to find |z|?

OpenStudy (phi):

\[ z-u = A e^{i \frac{\pi}{4}} \] if we plot that , we get the ray